OpenStudy (anonymous):

The lengths of the legs of a right triangle are 15 cm and 20 cm. What is the perimeter of the triangle? Note: You will need to use the Pythagorean theorem to find the hypotenuse. A. 35 cm B. 25 cm C. 60 cm D. 300 cm E. 20 cm

OpenStudy (anonymous):

is it 60???

OpenStudy (anonymous):

If a right triangle has legs of length A and B, and its hypotenuse is C, then: \[A^2 + B^2 = C^2\]

OpenStudy (anonymous):

I don't know thats why i need help

OpenStudy (anonymous):



OpenStudy (anonymous):

AB ^2 + BC ^2=AC ^2 15^2 +20 ^2 =AC^2 AC=sqrt(225+400)=sqrt(625)=25 cm hence perimeter=25+20+15=60cm